OpenStudy (anonymous):

Find all the roots. x^4-5x^2+4=0

OpenStudy (anonymous):

solve \[u^2-5u+4=0\] and then replace \(u\) by \(x^2\)

OpenStudy (anonymous):

\[u^2-5u+4=0\] \[(u-4)(u-1)=0\] \[u=4,u=1\] are the first steps

OpenStudy (anonymous):

then \(x^2=4\implies x=2\) or \(x=-2\) and \(x^2=1\implies x=1\) or \(x=-1\)

OpenStudy (anonymous):

@satellite73 Thanks :) That really helped!!